An elevator carrying 10 passengers through a height of 10 m from the ground has a mass of 2000 kg. Calculate the power required by the motor to lift the elevator if the time taken to lift the elevator to that height is 1 minute. 


1
Expert's answer
2022-02-13T12:12:59-0500

Given:

"m=2000\;\rm kg"

"h=10\;\rm m"

"t=60\:\rm s"


The power of the motor

"P=\frac{W}{t}=\frac{mgh}{t}=\frac{2000*9.8*10}{60}=3267\:\rm W"